when 20ml mixture of O2 and O3 is burned.volume became 29ml and invisible in basis payregalol solution.what is the volume% of oxygen gass in real mixture?
1
Expert's answer
2017-06-23T09:52:07-0400
The reaction equation
2O_3 → 3O_2
x+y=20
3/2 x+y=29
Thus,
1/2 x=29-20=9
So,
x=2(9)=18.
y=20-18=2.
The volume% of oxygen gas in real mixture is
2/20 100%=10%.